Hi, I was looking for help with setting up statistics problems in SPSS. I don't know how to enter the information correctly into SPSS to run the analysis. This is for a MANOVA analysis. I don't know how to input or code the data to enter it correctly into SPSS.60 3,973 2. A researcher is interested in the impact of a dietary intervention and/or exercise intervention on weight loss and cholesterol level. Study participants are randomly assigned to receive the dietary intervention and/or the exercise intervention. The data are given in the following table. Use MANOVA to evaluate the impact of the two interventions on the two outcomes, and determine whether there is interaction effect. an ID Diet Exercise Pounds Lost Cholesterol 0 96 126 166 168 173 178 190 194 0 4 0 is for a MANOVA and statement of results.
Answer #1

I have not SPSS but I do in R

y1=rep(0,8)
y2=rep(0,8)
y3=c(1,4,1,2,4,3,6,4)
y4=c(96,126,166,168,173,178,190,194)
fit=manova(cbind(y3,y4)~y4)
fit

manova(cbind(y3, y4) ~ y4)

Terms:
                      y4 Residuals
resp 1             5.572    15.303
resp 2          7885.875     0.000
Deg. of Freedom        1         6

Residual standard errors: 1.597012 7.536444e-15
Estimated effects may be unbalanced